Define and state the SI Unit of:
Electric field intensity
Calculate the magnitude of the electric field strength at a point 5 10 from an electron which is in a vacuum.
1
Expert's answer
2018-10-23T11:31:08-0400

E=k*|q|/r^2

r= 5*10^-10 m.

k=9*10^9 (constant value)

q=1,6*10^-19 (electron charge)

E=9*10^9*1,6*10^-19/(5*10^-10)^2=5,76*10^9 V/m.